# Correlated electronic structures and unconventional superconductivity in bilayer nickelate heterostructures

### v1

@@ Introduction
To set the stage for discussions of electronic bands on both DFT+$U$ and DFT+CDMFT levels, a realistic multi-orbital Hubbard model for the half-UC thin film is formulated, containing three parts: the non-interacting tight-binding (TB) model, the onsite interaction, and the chemical potential term
$H=\sum_{{\bf k\sigma}}\Psi_{{\bf k\sigma}}^{\dagger}H^{0}({\bf k})\Psi_{\bf k\sigma}+\sum_{i\tau}H_{i\tau}^{\mathrm{int}}-\mu\sum_{i\tau\gamma\sigma}n_{i\tau\gamma\sigma}.$ (1)
Here, the index $i$ enumerates the unit cells, $\tau$ the layer indices($A$ for the bottom layer, $B$ for the top layer), $\gamma$ the orbital indices ($x$ for $d_{x^{2}-y^{2}}$ orbital and $z$ for $d_{3z^{2}-r^{2}}$ orbital), $\sigma$ the spin states $\uparrow$ and $\downarrow$, $\mu$ the chemical potential, and $\Psi_{\bf k\sigma}=\left(d_{\bf kAx\sigma},d_{\bf kAz\sigma},d_{\bf kBx\sigma},d_{\bf kBz\sigma}\right)^{T}$. The chemical potential $\mu$ is used to control the electronic filling per Ni-layer per unit-cell $n=\sum_{i\tau\gamma\sigma}n_{i\tau\gamma\sigma}/N$ with $N$ the number of Ni sites in the system. We also introduce the inter-layer bonding (+) and anti-bonding (-) orbital basis $d_{z_{\pm}}=(d_{Az}\pm d_{Bz})/\sqrt{2}$ and $d_{x_{\pm}}=(d_{Ax}\pm d_{Bx})/\sqrt{2}$ as we will frequently use them.
@@
The low-energy bands’ TB model is obtained from the maximally-localized Wannier functions constructed by the Wannier90 package[36, 37]. As shown by the green dashed lines in Fig. 2b, the bands of Wannier TB align well the low-energy DFT bands. The TB model Hamiltonian in ${\bf k}$-space has the same matrix structure as in Ref. [11], reading
@@
Here, $T^{x/z}_{\bf k}$ ($T^{\prime x/z}_{\bf k}$) is the intra-layer (inter-layer) intra-orbital hopping, and $V_{{\bf k}}$ ($V_{{\bf k}}^{\prime}$) is the intra-layer (inter-layer) inter-orbital hopping. The onsite orbital level $\epsilon_{x/z}$ and hopping parameters are summarized in Table 2.
Compared to the TB model of the high-pressure bulk crystal, there are mainly two differences in the model parameters. The inter-layer intra-$z$-orbital hopping $t_{\perp,\text{0.5UC}}^{z}$=-0.439 eV is nearly 30% (0.2 eV) smaller in magnitude than its high-pressure bulk crystal counterpart ($t_{\perp,\text{HP}}^{z}$=-0.635 eV), while the crystal-field splitting $\Delta E_{\text{0.5\text{UC}}}=\epsilon_{x}-\epsilon_{z}$ between the $x$ and $z$ orbitals is enhanced to 0.519 eV, roughly $\sim$40 % larger than that $\Delta E_{\text{HP}}=0.367$ eV. These two factors greatly affect the bands and Fermi surface topology. Smaller $t_{\perp}^{z}$ means smaller bonding-anti-bonding splitting, which leads to the coexistence of $\delta$ electron pocket (located at $\Gamma$ point) formed by the $z_{-}$ band and the $\gamma$-hole pocket (located at $M$ point) formed by $z_{+}$ band, as shown in Fig. 2c-d for both hole-doped (with 0.2 hole doping per Ni) and un-doped scenarios. This coexistence, however, does not happen in both high-pressure [11] and ambient pressure bulk crystal[35].
The interacting part of Eq. 1 reads $H^{\mathrm{int}}=\sum_{i\tau}H^{\mathrm{int}}_{i\tau}$ with $H^{int}_{i\tau}$ the onsite multi-orbital interaction
@@
$H^{int}_{i\tau}$ contains the intra-orbital ($U$), inter-orbital ($U^{\prime}=U-2J$) Coulomb repulsions, and the Hund coupling $J$. The second line in Eq. 9 shows the spin-flip and pair-hopping terms. The realistic values of $U$ and $J$ are obtained by the constrained random phase approximation (cRPA)[38], which counts in the screening effect from the rest energy bands out of the low-energy $e_{g}$ subspace. The results obtained from cRPA are listed in Table 3. While there is slight anisotropy between $U_{x}$ and $U_{z}$, we take their average $\overline{U}=\sum_{\gamma}U_{\gamma}/2\approx$3.77 eV for both orbital and the Hund coupling J is taken as $J_{\mathrm{cRPA}}$=0.56 eV, which are slightly different from those of the high-pressure bulk crystal (3.79 and 0.61 eV)[17]. We also listed the inter-layer Coulomb repulsion $V$ ranging in1.01.3similar-toabsent1.01.3$\sim$ 1.0-1.3∼ 1.0 - 1.3 eV in the thin film, roughly 30% of $\overline{U}_{\mathrm{cRPA}}$. According to previous GW+EDMFT studies[17] on the high-pressure bulk crystal, $U$ can be further screened by the screening effect resulting from non-local $V$ but $J$ is un-screened. To mimic this screening effect and also systematically study the effect of $U$ on the correlated electronic structure, we tune the value of $U$ from 2 eV to $\overline{U}_{\mathrm{cRPA}}$=3.77 eV with fixed $J=J_{\mathrm{cRPA}}$=0.56 eV. We will show below that the Fermi surface topology will be substantially changed with $U$. The crystal structure, TB model obtained from wannier90 with all long-range hopping integrals and the cRPA interaction tensor $U_{ijkl}$ have been open-sourced in Zenodo[55].
CDMFT Studies
The main results of CDMFT are summarized in Fig.3, which in detail demonstrate how the physical quantities like quasi-particle weight, occupancy, effective energy level, spectral functions, and Fermi surfaces vary as one increases $U$. The calculations are performed at $T$=200 K and $n_{d}$=1.3 per site. More results on doping and temperature dependence can be found in S.M.
As one increases $U$, we observe an increase in effective energy splitting $\epsilon^{\text{eff}}_{z_{-}}-\epsilon^{\text{eff}}_{z_{+}}$ between the $z_{+}$ and $z_{-}$ orbitals as $U$ increases, where $\epsilon^{\text{eff}}_{\alpha}=\epsilon_{\alpha}+\mathrm{Re}\Sigma{\alpha(i0^{+})}-\mu$. Thus the effective energy level $\epsilon^{\text{eff}}_{z_{-}}$ ($\epsilon^{\text{eff}}_{z_{+}}$) of the $z_{-}$ ($z_{+}$ ) orbital is pushed upward (downward) as seen in Fig. 3a. For $U$<3.25 eV, $\epsilon^{\text{eff}}_{z_{\pm}}$ deviates slowly from its bare value $\epsilon_{z_{\pm}}$ (dashed lines). The deviation increases rapidly when $U\geq 3.5$ eV, as indicated by the arrows. This is consistent with the picture of strong inter-layer electronic correlations between the $z$ orbitals [25], which becomes more prominent as $U$ increases. On the contrary, $\epsilon^{\text{eff}}$ for the $x$ orbitals remains gradually increasing even for larger $U$. $U=3.5$ eV could be regarded as a characteristic interaction strength where the correlation effect becomes significant.
The $U$-dependent orbital occupancy is given in Fig. 3b. There is a self-doping effect with the electrons transfer from the $x$ orbitals to the $z$ orbitals (note that $n_{d}$=1.3 fixed). As $U$ is tuned from 2 to 3.77 eV, the total filling in the $z$ orbitals increases from $\sim$0.76 to $\sim$0.9 while the total filling in the $x$ orbitals decreases from $\sim$0.54 to $\sim$ 0.4. The quasi-particle weight $Z$ of the $x_{\pm}$ orbitals ranges from $\sim$0.7 to $\sim$0.5 as $U$ is tuned down, which results from a small filling per spin ($\sim$ 0.3 to $\sim$0.2). The $Z$ of $z_{\pm}$ orbitals decreases from $\sim$0.45 to $\sim 0.2$ for the $z_{+}$ orbital and to $\sim$ 0.08 for $z_{-}$ orbital, showing that $z$ orbitals are much more correlated than the $x$ orbitals. The reason for this is that the self-doping effect from $x$ to $z$ orbital makes the filling in each $z$ orbital approach to half filling per spin (from $\sim$0.38 at $U$=2 to $\sim$0.45 at $U=3.77$.)
Associated with the enhanced effective level splitting as $U$ increases, the ${\bf k}$ spectral functions $A({\bf k},\omega)$ and Fermi surface evolve correspondingly. From panel 3e-h, one can see that both $\delta$-band at $\Gamma$ point and $\gamma$-band at $M$-point at low-energy become more flat and gradually fade away (being less coherent) from the Fermi energy. First, as mentioned before, there is a downward or upward shift of $\epsilon^{\text{eff}}_{z_{\pm}}$. Second, the quasi-particle spectral weight (see Fig.3b) decreases, which makes especially the $\gamma$ band more flat near the Fermi energy. Furthermore, the scattering rate is enhanced when $U$ increases (not shown). As a result, the Fermi surface of $\delta$ electron pocket gradually disappears, as seen from Fig.3i-l . Simultaneously, the $\gamma$ hole pocket also shrinks in size and becomes less coherent (as evidenced by its decreased intensity in colormap). At $U=3.77$, the $\gamma$ band becomes flat near $M$ point, and its band-top touches the Fermi energy, which gives rise to the Fermi surface shown in Fig. 3n.
Pairing symmetry and mechanism (CDMFT+RPA)
The superconducting pairing symmetry has been studied with a modified RPA approach, where the quasiparticle Hamiltonian from the CDMFT is taken as an input (see Method and SM.4 for details). We have assumed that effective interaction used in RPA is in the same form as the one in Eq. 9 and we fix $J_{\textrm{eff}}=U_{\textrm{eff}}/6$ in the calculations. The effective interaction strength $U_{\textrm{eff}}$ is a tuning parameter. It could be regarded as the strength of some kinds of residual interactions renormalized by a factor $Z^{2}$, where $Z$ is the quasi-particle weight. Since $Z$ of $d_{z^{2}}$ orbital is quite small ($\sim 0.2$ for $U=3.5$ eV), the effective interaction strength $U_{\textrm{eff}}$ should also be very small.
The RPA results for $U=3.5$ eV and $n=1.3$ are summarized in Fig.  4. In Fig. 4a, we show the typical RPA-renormalized spin susceptibility. The dominant wavevector $\mathbf{Q}_{1}$ corresponds to the nesting between the $\beta$-pocket and $\gamma$-pocket as shown in Fig. 4c. In FIG. 4b, we show the dependence of effective pairing strength[39] $\lambda$ on $U_{\textrm{eff}}$ for various pairing symmetries. It is clear that larger $U_{\textrm{eff}}$ leads to stronger superconducting instability. The leading pairing symmetry is always a $s^{\pm}$-wave, where the gap function on the Fermi surface is depicted in Fig. 4c. The coincidence of $\mathbf{Q}_{1}$ in FIG. 4a and the one connects the largest gap in magnitude in c supports the spin-fluctuations-mediated pairing mechanism. In real space, the strongest pairing is between the inter-layer $d_{z^{2}}$ orbitals as shown in Fig. 4d, which is consistent with the fact that the Fermi surface patches connected by $\mathbf{Q}_{1}$ on $\beta$ and $\gamma$ bands shown in Fig. 4c are mainly formed by $d_{z^{2}}$ orbitals. We also perform similar calculations for $3\leq U\leq 3.77$, and $s^{\pm}$-wave pairing is always dominant (see SM.4 for more details). All in all, the most possible pairing symmetry is $s^{\pm}$-wave pairing for realistic parameters.
Discussion
At first, we would like to discuss to what extent the correlation effect is in the 327 films. The on-site Coulomb repulsion from cRPA is $U\sim 3.77$ eV, consistent with the typical value of $U=3-4$ eV for the transition metals. Due to the strong interlayer coupling, the multi-orbital tight-binding model involves four bands with a total bandwidth around 4 eV, which is comparable to the on-site Hubbard $U$. The situation is similar to the one in the iron-based superconductor[40, 41, 42, 43], suggesting that the nickelate 327 systems also lie in an intermediate correlation regime.
Though the correlation is not too strong, its effect is non-negligible. We have performed a calculation based on fluctuation-exchange (FLEX) approximation[44, 45, 29] at the same filling $n=1.3$. As a weak coupling approach, our FLEX calculations a restricted to $U\leq$1.8 eV, above which it may break down in convergence. In the weak $U$ region, the FLEX can reproduce the Fermi surface obtained in CDMFT for $U\leq$ 3.0 (see SM.5). However, FLEX fails to sees an Fermi surface topology similar to Fig. 3k obtained in CDMFT at $U=3.5$ eV. This indicates that FLEX can not describe the strong inter-layer correlation which may play an important role in the electronic structure of the 327 films.
Then we discuss the relation between strained thin films and pressurized bulks. The coupling between interlayer $d_{z^{2}}$ orbitals ($t_{\perp}^{z}/J_{\perp}^{z}$) are believed to be crucial in the pressurized bulk sample[13, 29, 19]. However, thin films have a much longer interlayer distance than the one in pressurized bulk sample, which leads to much weaker inter-bilayer couplings, say about $30\%$ lower for $t^{z}_{\perp}$ and $50\%$ lower for $J^{z}_{\perp}$ (since it proportion to $(t^{z}_{\perp})^{2}$). It would be a challenge for the theories strongly depends on $t_{\perp,z}$ and/or $J^{z}_{\perp}$ to explain the comparable on-set Tc of the two systems. On the contrary, the resultant Fermi surface from CDMFT by including the correlation effect is similar to the one in high pressure, except that the $\delta$-pocket is much closer to the Fermi surface in the film case. So, it would be much easier to understand why the thin films has very similar onset $T_{\mathrm{C}}$ with a much smaller $t^{z}_{\perp}$ given that the correlated effect has been included correctly.
In conclusion, based on structural analysis of superconducting bilayer nickelate thin films, combining STEM and ab initio calculations, a realistic multi-orbital Hubbard model is constructed. Considering a finite screening effect, in the regime lower than the cRPA calculated interaction parameter, Fermi surface topology is qualitatively altered by the Coulomb repulsive $U$, featuring an indispensable role of the electron correlations. The modified RPA in this regime gives rise to a predominant $s^{\pm}$-wave pairing instability driven by spin-fluctuations arising from Fermi surface nesting between $d_{z^{2}}$-orbital bands. With intermediate correlation correctly included to construct the electronic structures of the superconducting thin-film bilayer nickelates, the similar order-of-magnitude onset $T_{\mathrm{C}}$ between pressured bulks and thin films may imply a critical yet secondary influence of the lowered $d_{z^{2}}$-orbital interlayer coupling in the strained films. These results set a foundation for the design of higher $T_{\mathrm{C}}$ systems within the bilayer nickelate family.
{methods}
STEM experiment All electron microscopy analyses were conducted using a FEI Themis Z transmission electron microscope operated at 200 kV. This microscope is equipped with a Cs Probe Corrector (DCOR) and a high-brightness field-emission gun (X-FEG) with a monochromator to enhance resolution and contrast. For scanning transmission electron microscopy (STEM) high-angle annular dark field (HAADF) and annular bright field (ABF) imaging, the inner and outer acquisition angles ($\beta$1 and $\beta$2) were set to 90 and 200 mrad for HAADF images, and 12 and 23 mrad for ABF images, respectively. The convergence angle was maintained at 25 mrad. The beam current was adjusted to approximately 40 pA for HAADF and ABF imaging. Cross-sectional STEM specimens were prepared using a FEI Helios G4 HX dual-beam focused ion beam/scanning electron microscope (FIB/SEM) system. The simulation of ABF images was performed using the multi-slice method implemented in the QSTEM software. The simulation parameters were carefully set to match those used during the electron microscopy characterization, ensuring that the simulated images accurately reflect the experimental conditions.
@@ Discussion
CDMFT calculation We apply the CDMFT[46, 47] method to study the correlated electronic structure of the half-UC thin film. In the bonding and anti-bonding basis of ${z_{+}}$ ${z_{-}}$ ${x_{+}}$ ${x_{-}}$, the cluster impurity problem is equivalent to an effective four-orbital single impurity model with diagonal crystal field splitting among these orbitals. The numerically exact hybridization-expansion continuous-time quantum Monte Carlo (CT-HYB)[48, 49] is adopted to solve the quantum impurity model. To calculate the momentum-resolved spectral function, we adopt the maximum-entropy analytic continuation method to continue the self-energy from the Matsubara frequency axis to the real frequency axis[50].
@@
CDMFT+RPA calculation The CDMFT+RPA method is a modified RPA approach that the infinite sum of bubble diagrams, i.e. the random phase approximation (RPA), is performed upon the quasi-particle Hamiltonian from CDMFT. We assume that the effective interaction has the same form as bare interaction only with strength renormalized by the quasi-particle weight. The advantage of CDMFT+RPA method is that small effective interaction actually corresponds to intermediate bare interaction. The RPA-renormalized spin and charge susceptibilities can investigate spin-density wave (SDW), charge-density wave (CDW), and pairing instabilities. There exist critical interaction strength, above which RPA results become unreliable. Near critical values, enhanced spin-fluctuations can mediate an effective attraction between quasi-particles close to the Fermi surface, and serve as glues for the formation of Cooper pairs. The effective attraction is treated by mean-field theory to derive the self-consistent superconducting gap equation. Near $T_{\mathrm{C}}$, the gap equation can be linearized, and becomes a standard eigenvalue problem. The largest eigenvalue of the effective pairing interaction vertex matrix determines $T_{\mathrm{C}}$, and corresponding eigenvectors give the distribution of gap functions, from which we can determine the pairing symmetry[39].
@@
0.1 SM1. Additional structural information
The lattice structure parameters used in the calculations were taken from a superconducting La2.85Pr0.15Ni2O7 film and a non-superconducting La3Ni2O7 film grown on SrLaAlO4 substrates[34]. Fig. S1 shows the annular bright-field (ABF) images of the cross-section of La2.85Pr0.15Ni2O7 sample, projected along [100]pc, [010]pc and [110]pc. We analyzed the positions of oxygen atoms in these three projections and selected part of the data to show in Fig. 1, which reflects that the La2.85Pr0.15Ni2O7 thin film has almost no oxygen octahedral rotation under compressive epitaxial strain. The atomic spacing used for modeling is obtained by measuring directly on High-angle Annular Dark Field (HAADF) images with calibration from XRD lattice constants. Fig. S2 identifies several key atomic spacing values of La2.85Pr0.15Ni2O7 and La3Ni2O7 films, which are also listed in Tables 1 and Tables S1. At the same time, considering the error of the electron microscope in the measurement, the listed values show only two significant digits after the decimal point.
0.2 SM2. Ab initio Calculations
@@
0.3 SM3. Doping and Temperature Dependence of CDMFT results
Fig. S5 shows the doping dependence of $A({\bf k},\omega)$ and FS at $U=3.5$ eV and indicated fillings. As one decreases the filling $n=1.5$ to $n=1.1$ (equivalent to increasing hole doping from 0 hole to 0.4 hole per Ni-$e_{g}$ orbitals), one can see that the $\gamma$ ($\delta$) band moves upward (downward), which can be explained by a reduced effective level splitting between $z_{+}$ and $z_{-}$ orbitals. The Fermi surface at $n=1.1$ (Fig. S5f) resembles the one seen in DFT (Fig. 2c) at $n=1.33$. The reason is that the inter-orbital correlation strength decreases when the filling per orbital is reduced. As a result, the $\gamma$ hole ($\delta$ electron) pocket gradually appears and expands as $n$ reduces.
Fig. S6 shows the temperature dependence of $A({\bf k},\omega)$ and FS, which changes only qualitatively as one decreases of temperature from $T$=200 K to $T$=50 K. The main difference is that the low-energy electrons become more coherent at lower $T$, leading to sharper low-energy bands and sharper Fermi surfaces.
0.4 SM4. CDMFT+RPA
CDMFT+RPA is based on the quasi-particle Hamiltonian from CDMFT. As shown by our CDMFT calculation, the Fermi surface topology can be greatly affected at larger $U\geq$>3.5 eV. As the pairing is mainly controlled by the low-energy excitations, a better starting point for RPA is to use the quasi-particle Hamiltonian obtained from CDMFT. We call such treatment CDMFT+RPA. The quasi-particle Hamiltonian $\tilde{H}_{{\bf k}}^{\mathrm{QP}}$ is obtained from the quasi-particle approximation of self-energy
@@
FIG. S7 illustrates that the correspondence between $\mathbf{Q}_{1}$ in the distribution of RPA-renormalized spin susceptibility and the FS nesting still keeps for different parameters. As the shape of FS changes for different fillings and bare interactions in DMFT, the wave vector $\mathbf{Q}_{1}$ also varies accordingly. FIG. S7d-f demonstrate that the gap functions on FS patches connected by $\mathbf{Q}_{1}$ possess opposite signs, which is similar to the $d_{x^{2}-y^{2}}$-wave pairing in the cuprates and the $s^{\pm}$-wave pairing in the iron-based superconductors.
@@
FIG. S8 presents the dependence of $\lambda$ on $U_{\textrm{eff}}$ for $U=3.77$ eV and $U=3$ eV. The leading pairing symmetry is always the $s^{\pm}$ wave pairing before RPA-renormalized spin susceptibility diverges. For small bare interaction in FIG. S8b, there is obvious competition between $s^{\pm}$-wave and $d_{xy}$-wave pairings. With stronger bare interaction, the $s^{\pm}$-wave pairing instability is leading only as shown in FIG. S8a. Both results further support the main conclusion that the most possible pairing symmetry is $s^{\pm}$-wave pairing for realistic parameters in the main text.
0.5 SM5. FS in Fluctuation-exchange approximation
@@
Fluctuation-exchange (FLEX) approximation[44] have been developed to study the pairing instability of Hubbard models. FLEX is by nature a weak-coupling approach and it breakdowns when $U$ is stronger. Compared with normal RPA, FLEX can reach slightly higher correlation strength as the self-energy is incorporated. Here, we apply the FLEX calculation to the same multi-orbital Hubbard model of the ultra-thin film using a modified FLEX package developed by Witt et al.[45]. The Fermi surfaces at filling $n=1.3$ are shown in Fig. S9a-b. It turns out the $\delta$-pocket remains at $\Gamma$ point when $U$ is increased from $U=1$ to $U=1.8$ above which FLEX breaks down. While the FS of FLEX at the same filling $n=1.3$ is similar to that given by CDMFT at $U\leq 3$ (Fig. 3i-j), FLEX fails to reproduce FS as in Fig. 3l in the safe radius of $U$. Note that the largest $U$ that FLEX can reach depends on the filling and density of states. We confirm that our FLEX calculation can reach $U\leq 3$ for the TB model of high-pressure bulk crystal, which has a less flat $\gamma$-band than the ultra-thin film.
Figure S1: | Annular bright-field (ABF) images with a larger fields of view of the cross-section of La2.85Pr0.15Ni2O7 film. a-c, The ABF images, projected along [100]p, [010]pc and [110]pc, respectively.
Figure S2: |High angle Annular Dark Field (HAADF) images with atomic spacings indicated. a, Atomic spacing of the La2.85Pr0.15Ni2O7 film on SrLaAlO4 substrate. b, Atomic spacing of the La3Ni2O7 film on SrLaAlO4 substrate. The atom spacing in the substrate and the film was calibrated using the corresponding XRD values.
@@
Table S1: | The lattice parameters of La3Ni2O7 film and substrate SrLaAlO4. STEM measurements are calibrated by the lattice constants obtained from X-ray diffraction experiments. The term "La-La intra" refers to the distance between La atoms in the c-axis direction in the perovskite layer. The term "La-La inter" refers to the distance between adjacent La atoms in the c-axis direction between the perovskite bilayers. All lattice parameters are represented numerically using pseudocube cells.

### v2

@@ Discussion
Ab initio multi-orbital Hubbard model. We construct the realistic multi-orbital Hubbard model for the half-UC thin film, including the non-interacting tight-binding (TB) model and the onsite interaction term
$H=\sum_{{\bf k\sigma}}\Psi_{{\bf k\sigma}}^{\dagger}H^{0}({\bf k})\Psi_{\bf k\sigma}+\sum_{i\tau}H_{i\tau}^{\mathrm{int}}.$ (1)
We use $i$ to enumerate the unit cells, $\tau$ the layer indices ($A$ for the bottom layer, $B$ for the top layer), $\gamma$ the orbital indices ($x$ for $d_{x^{2}-y^{2}}$ orbital and $z$ for $d_{z^{2}}$ orbital), $\sigma$ the spin states $\uparrow$ and $\downarrow$, and $\Psi_{\bf k\sigma}=\left(d_{\bf kAx\sigma},d_{\bf kAz\sigma},d_{\bf kBx\sigma},d_{\bf kBz\sigma}\right)^{T}$.
@@
The low-energy bands’ TB model is obtained from the maximally localized Wannier functions constructed by the Wannier90 package[46, 47]. The TB model Hamiltonian in ${\bf k}$-space has the same matrix structure as in Ref. [11], reading
@@
Here, $T^{x/z}_{\bf k}$ ($T^{\prime x/z}_{\bf k}$) is the intra-layer (inter-layer) intra-orbital hopping, and $V_{{\bf k}}$ ($V_{{\bf k}}^{\prime}$) is the intra-layer (inter-layer) inter-orbital hopping. The hopping parameters are summarized in Table 2.
The interacting part of Eq. 1 reads $H^{\mathrm{int}}=\sum_{i\tau}H^{\mathrm{int}}_{i\tau}$ with $H^{int}_{i\tau}$ the Kanamori type two-orbital onsite interaction
@@
$H^{int}_{i\tau}$ contains the intra-orbital ($U$), inter-orbital ($U^{\prime}=U-2J$) Coulomb repulsions, and the Hund coupling $J$. The second line in Eq. 9 shows the spin-flip and pair-hopping terms. The realistic values of $U$ and $J$ are obtained by the constrained random phase approximation (cRPA)[38], which counts in the screening effect from the rest energy bands out of the low-energy $e_{g}$ subspace. The results obtained from cRPA are listed in Table 3. Although there is slight anisotropy between $U_{x}$ and $U_{z}$, we take their average $\overline{U}=\sum_{\gamma}U_{\gamma}/2\approx$3.77 eV for both orbital and the Hund coupling J as $J_{\mathrm{cRPA}}$ = 0.56 eV, which is slightly different from those of the high pressure bulk crystal (3.79 and 0.61 eV)[17]. We also listed the inter-layer Coulomb repulsion $V$ ranging in1.01.3similar-toabsent1.01.3$\sim$ 1.0-1.3∼ 1.0 - 1.3 eV in the thin film, roughly 30% of $\overline{U}_{\mathrm{cRPA}}$. The crystal structure, TB model obtained from wannier90 with all long-range hopping integrals and the cRPA interaction tensor $U_{ijkl}$ have been open-sourced in Zenodo[58].
@@
CDMFT calculation. We apply the CDMFT[49, 50] method to study the correlated electronic structure of the half-UC thin film. In the bonding and anti-bonding basis of ${z_{+}}$ ${z_{-}}$ ${x_{+}}$ ${x_{-}}$, the cluster impurity problem is equivalent to an effective four-orbital single impurity model with diagonal crystal field splitting among these orbitals. The numerically exact hybridization-expansion continuous-time quantum Monte Carlo (CT-HYB)[51, 52] is adopted to solve the quantum impurity model. To calculate the momentum-resolved spectral function, we adopt the maximum-entropy analytic continuation method to continue the self-engery from the Matsubara frequency axis to the real frequency axis[53].
@@
CDMFT+RPA calculation. The CDMFT+RPA method is a modified RPA approach in which the infinite sum of bubble diagrams, i.e., the random phase approximation (RPA), is performed upon the quasi-particle Hamiltonian from CDMFT. We assume that the effective interaction has the same form as bare interaction with strength only renormalized by the quasi-particle weight. The advantage of the CDMFT+RPA method is that small effective interaction actually corresponds to intermediate bare interaction. The RPA-renormalized spin and charge susceptibilities can investigate spin-density wave (SDW), charge-density wave (CDW), and pairing instabilities. There exists critical interaction strength, above which RPA results become unreliable. Near critical values, enhanced spin fluctuations can mediate an effective attraction between quasi-particles close to the Fermi surface and serve as glues for forming Cooper pairs. The effective attraction is treated by mean-field theory to derive the self-consistent SC gap equation. Near SC critical temperature $T_{c}$, the gap equation can be linearized and becomes a standard eigenvalue problem. The largest eigenvalue of the effective pairing interaction vertex matrix determines $T_{c}$, and corresponding eigenvectors give the distribution of gap functions, from which we can determine the pairing symmetry[39].
@@
SM1. Additional structural information
The lattice structure parameters were taken from a superconducting La2.85Pr0.15Ni2O7 film and a non-superconducting La3Ni2O7 film grown on SrLaAlO4 substrates[34]. Fig. S1 shows the annular bright-field (ABF) images of the cross-section of La2.85Pr0.15Ni2O7 sample, projected along [100]pc, [010]pc and [110]pc. We analyzed the positions of oxygen atoms in these three projections and selected part of the data to show in Fig. 1, which reflects that the La2.85Pr0.15Ni2O7 thin film has almost no oxygen octahedral rotation under compressive epitaxial strain. The atomic spacing used for modeling is obtained by measuring directly on High-angle Annular Dark Field (HAADF) images with calibration from XRD lattice constants. Fig. S2 identifies several key atomic spacing values of La2.85Pr0.15Ni2O7 and La3Ni2O7 films, which are also listed in Tables 1 and Tables S1. At the same time, considering the error of the electron microscope in the measurement, the listed values show only two significant digits after the decimal point.
SM2. Ab initio Calculations
@@
SM3. Doping and Temperature Dependence of CDMFT results
Fig. S5 shows the doping dependence of $A({\bf k},\omega)$ and FS at $U=3.6$ eV and indicated fillings. As one decreases the filling $n=1.4$ to $n=1.2$ (equivalent to increasing hole doping from 0.1 hole to 0.3 hole per Ni-$e_{g}$ orbitals), one can see that the $\gamma$ ($\delta$) band moves upward (downward), which can be explained by a reduced effective level splitting between $z_{+}$ and $z_{-}$ orbitals. The Fermi surface at $n=1.2$ (Fig. S5f) resembles the one seen in DFT (Fig. 2c) at $n=1.33$. The reason is that the inter-orbital correlation strength decreases when the filling per orbital is reduced. As a result, the $\gamma$ hole ($\delta$ electron) pocket gradually appears and expands as $n$ reduces.
Fig. S6 shows the temperature dependence of $A({\bf k},\omega)$ and FS at $U=3.6$, which changes only qualitatively as one decreases of temperature from $T$=200 K to $T$=50 K. The main difference is that the low-energy electrons become more coherent at lower $T$, leading to sharper low-energy bands and sharper Fermi surfaces.
SM4. CDMFT+RPA
CDMFT+RPA is based on the quasi-particle Hamiltonian from CDMFT . As shown by our CDMFT calculation, the Fermi surface topology can be greatly affected at larger $U$>3 eV. As the pairing is mainly controlled by the low-energy excitations, a better start point for RPA is to use the quasi-particle Hamiltonian obtained from CDMFT. We call such treatment CDMFT+RPA. The quasi-particle Hamiltonian $\tilde{H}_{{\bf k}}^{\mathrm{QP}}$ is obtained from the quasi-particle approximation of self-energy
@@
FIG. S7 illustrates that the correspondence between $\mathbf{Q}_{1}$ in the distribution of RPA-renormalized spin susceptibility and FS nesting still holds for different parameters. As the shape of FS changes for different fillings and bare interactions in DMFT, the wave vector $\mathbf{Q}_{1}$ also varies accordingly. FIG. S7b,d,f,h demonstrate that the gap functions on FS patches connected by $\mathbf{Q}_{1}$ possess opposite signs, which is similiar to the $d_{x^{2}-y^{2}}$-wave pairing in the cuprates and the $s^{\pm}$-wave pairing in the iron-based superconductors.
@@
FIG. S8 presents the dependence of $\lambda$ on $U_{\textrm{eff}}$ for $U=3.77$ eV, $U=3.6$ eV, and $U=3$ eV. The leading pairing symmetry is always the $s^{\pm}$ wave pairing before RPA-renormalized spin susceptibility diverges. For small bare interaction in FIG. S8c, there is obvious competition between $s^{\pm}$-wave and $d_{xy}$-wave pairings. With stronger bare interaction, only the $s^{\pm}$-wave pairing instability is leading as shown in FIG. S8a. Both results further support the main conclusion that the most possible pairing symmetry is $s^{\pm}$-wave pairing for realistic parameters in the main text.
SM5. FS in Fluctuation-exchange approximation
@@
Fluctuation-exchange (FLEX) approximation[44] has been developed to study the pairing instability of Hubbard models. FLEX is by nature a weak-coupling approach and it breakdowns when $U$ is stronger. Compared with normal RPA, FLEX can reach slightly higher correlation strength as the self-energy is incorporated. Here, we apply the FLEX calculation to the same multi-orbital Hubbard model of the ultra-thin film using a modified FLEX package developed by Witt et al.[45]. The Fermi surfaces at filling $n=1.3$ are shown in Fig. S9a-b. It turns out the $\delta$-pocket remains at $\Gamma$ point when $U$ is increased from $U=1$ to $U=1.8$ above which FLEX breaks down. While the FS of FLEX at the same filling $n=1.3$ is similar to that given by CDMFT at $U\leq 3$ (Fig. 3i-j), FLEX fails to reproduce FS as in Fig. 3l in the safe radius of $U$. Note that the largest $U$ that FLEX can reach depends on the filling and density of states. We confirm that our FLEX calculation can reach $U\leq 3$ for the TB model of high-pressure bulk crystal, which has a less flat $\gamma$-band than the ultra-thin film.
@@
Table S1: | The lattice parameters of La3Ni2O7 film and substrate SrLaAlO4. STEM measurements are calibrated by the lattice constants obtained from XRD experiments. The term "La-La intra" refers to the distance between La atoms in the c-axis direction in the perovskite layer. The term "La-La inter" refers to the distance between adjacent La atoms in the c-axis direction between the perovskite bilayers. The "cif" column lists the parameters from standard ".cif" file. All lattice parameters are represented numerically using pseudocube cells.
@@
Figure S1: | Annular bright-field (ABF) images of the cross-section of La2.85Pr0.15Ni2O7 film. a-c, The large-scale ABF images, projected along [100]p, [010]pc and [110]pc, respectively, delineate the boundaries of the film and the substrate.
Figure S2: |High angle Annular Dark Field (HAADF) images with atomic spacings indicated. a, Atomic spacing of La2.85Pr0.15Ni2O7 film on SrLaAlO4 substrate. b, Atomic spacing of the La3Ni2O7 film on SrLaAlO4 substrate. The atom spacings in the substrates and the films are calibrated using the XRD measurements.